OpenStudy (uri):

Find the point(s) of intersection of the parabola with equation y = x2 - 5x + 4 and the line with equation y = 2x - 2

OpenStudy (mathteacher1729):

Note how both of these are equal to "y" y = x^2 - 5x + 4 y = 2x - 2 or x^2 - 5x + 4 = y = 2x - 2 now just "take out the middle man" (the y) and solve x^2 - 5x + 4 = 2x - 2 Can you do this?

OpenStudy (uri):

from quadratic equation?

OpenStudy (mathteacher1729):

First step is to move all the variables to one side of the equal sign.

OpenStudy (uri):

ikr?

OpenStudy (mathteacher1729):

x^2 - 5x + 4 = 2x - 2 add 2 to both sides x^2 - 5x + 6 = 2x subtract 2x from both sides x^2 - 7x + 6 = 0 See if you can factor the left side (if not, use quadratic formula)

OpenStudy (uri):

Thankss :D